OpenStudy (anonymous):

Find the sum to infinity of the series 40 - 24 + 14.4 -....

OpenStudy (anonymous):

What do you mean by ...... on your question?

Parth (parthkohli):

Broaden it please

OpenStudy (anonymous):

Seems like the values are getting smaller so it has to be smaller than 40

OpenStudy (anonymous):

..... means the rest of the terms in the series i think its a geometric??

OpenStudy (anonymous):

Yeah I really don't know how the series changes as n= goes to infinity.

OpenStudy (asnaseer):

this is a geometric series where you multiply by \(\displaystyle-\frac{3}{5}\) each time

OpenStudy (asnaseer):

use the formulae described here for the sum to infinity: http://en.wikipedia.org/wiki/Geometric_progression

OpenStudy (anonymous):

its n = 1 to infinity, i assume

OpenStudy (anonymous):

no 0 to infinity

OpenStudy (asnaseer):

\[\sum_{k=0}^{\infty}ar^k=\frac{a}{1-r}\]

OpenStudy (anonymous):

ok - i checked your link its a / 1 - r

OpenStudy (anonymous):

thanx

OpenStudy (asnaseer):

in your case a=40 and r=-3/5

OpenStudy (anonymous):

right 40 / 1 + 0.6 = 25

OpenStudy (asnaseer):

correct